OpenStudy (anonymous):

-cos135+isin135/sin105-icos105=?

OpenStudy (solomonzelman):

\[-Cos(A)=Cos(-A)\]

OpenStudy (solomonzelman):

sin(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B cos(A+B)=cos A cos B - sin A sin B cos(A-B)=cos A cos B + sin A sin B try braking these cosines into exact numbers. For example: \[-Cos(135)=Cos(-135)=Cos(-45-90)\]

OpenStudy (solomonzelman):

Do you know how to expand this?

OpenStudy (solomonzelman):

Alright....

OpenStudy (anonymous):

how to cis?

OpenStudy (solomonzelman):

\[Cos(-45-90)=Cos(-45)~~Cos(90)~~-~~Sin(-45)Sin(90)\] \[-Cos(45)~~Cos(90)~~+~~-Sin(45)Sin(90)\]\[-Cos(45)~~Cos(90)~~-~~Sin(45)Sin(90)\]

OpenStudy (solomonzelman):

you can simplify this before anything Cos45=Sin45 and Cos(90)=0

OpenStudy (solomonzelman):

And sin(90)=1

OpenStudy (solomonzelman):

\[-Sin(45)~-Sin(45)\]\[-2~Sin(45)\]\[Sin(45)= \sqrt{2}/2\]So,\[-2 \times( \sqrt{2}/2)\]\[-\sqrt{2}\]So\[-Cos(135)= -\sqrt{2}\]

OpenStudy (solomonzelman):

\[-\cos135+isin135/\sin105-icos105\]Now we have \[-\sqrt{2}+isin135/\sin105-icos105\]

OpenStudy (solomonzelman):

Did you mean i sin(135) ?

OpenStudy (anonymous):

yeap thanks so much